Calculate the gross and net calorific values of coal having the following



compositions, carbon = 85%, hydrogen = 8%, sulphur = 1%, nitrogen = 2%,



ash = 4%, latent heat of steam = 587 cal/gm

1
Expert's answer
2022-01-18T13:41:03-0500

Solution:

Dulong's formula for calculating the calorific value can be used:

Gross calorific value (GCV) = (1/100) × [8080C + 34500(H - O/8) + 2240S],

where C, S, O, H are the percentages of carbon, sulphur, oxygen and hydrogen.

So, substitute all the given values in the formula and calculate so that you will get GCV:

GCV = (1/100) × [8080×85 + 34500(8 - 0/8) + 2240×1] = 9650.4 (kcal/kg)

Gross calorific value is 9650.4 kcal/kg


Apply the formula for net calorific value (NCV): NCV = (GCV - 0.09H×587)

Therefore,

NCV = (9650.4 - 0.09×8×587) = 9227.76 (kcal/kg)

Net calorific value is 9227.76 kcal/kg


Answer:

The gross calorific value (GCV) of coal is 9650.4 kcal/kg

The net calorific value (NCV) of coal is 9227.76 kcal/kg